OpenStudy (anonymous):

Using point slope form or slope intercept form solve the following. Find an equation with slope 0 and containing points (4,2)

OpenStudy (anonymous):

y-y0 = m*(x-x0) => y-2=0(x-4) => y = 2 should be the answer.

OpenStudy (anonymous):

i got that but.. it doesnt contain points (4,2)

OpenStudy (anonymous):

if you graph it.

OpenStudy (anonymous):

in effect the equation is y-2 = 0(x-4), when you solve it you get y=2

OpenStudy (anonymous):

so then it doesn't contain points (4,2) because it does contain the y coordinate to be 2 but not the x coordinate.

OpenStudy (anonymous):

y=